I need help to find the difference in coverting
a) 1 mile and 1 kilometer =_________km
b) 3 miles and 5000 meters =_________ft
c)120 yards and 110 meters _____ft
d) 1 quart and 1 liter =______qt
e) 10 pounds and 5 kilograms =______lb
can you please help me I confused about the convert.

Answer by ewatrrr(24785)

 
Hi,
Might recommend finding a calculator that allows You to do Unit conversions
1mile = 1.609344km
1mile = 1609.344 meters
1mile = 5280ft and 3mi b) 3*5280ft 5000meter = 5000*3.2808ft
1 meter = 3.2808ft
1yd = 3ft 120yd c) =120*3ft = 360ft 110meters = 110*3.2808ft
1L = 1.0567Qt(US)
1 kilogram = 2.2046lb 5 kilogram e) 5* 2.2046lb = 11.0231lb